# Equivalence of Families of Polycyclic Codes over Finite Fields

### v1

@@ 2 Preliminaries
    The set $\varphi(C)$ is an ideal of the polynomial ring $\mathbb{F}_{q}[x]/\langle x^{n}-\vec{a}(x)\rangle,$ with $\vec{a}(x)=a_{n-1}x^{n-1}+\ldots+a_{1}x+a_{0}.$
@@
    There is a monic polynomial $g(x)\in\mathbb{F}_{q}[x]$ of least degree which divides $x^{n}-\vec{a}(x)$ and $\varphi(C)=\langle g(x)\rangle.$
@@
    The set $\{g(x),xg(x),\ldots,x^{{n-\deg(g)-1}}g(x)\}$ forms a basis of $\varphi(C)$ and the dimension of $C$ is $n-\deg(g(x)).$
@@
Let $C\subseteq\mathbb{F}_{q}^{n}$ be a non-zero (right) polycyclic code of length $n$ over $\mathbb{F}_{q}$. Then the standard generator of $C$ is the monic polynomial $g(x)$ of least degree in $\mathbb{F}_{q}[x]/\langle f(x)\rangle$ such that $\varphi(C)=\langle g(x)\rangle$. In this paper, we refer to the standard generator of $C$ as ”the generator”.
@@
[Redei1950, Lemma 1] Let $m,n\geq 1$ be integers and $a,b\in\mathbb{F}_{q}^{*}$. Then $\gcd\left(x^{n}-a,x^{m}-b\right)$ has degree 0 or $d:=\gcd(m,n)$ (over an arbitrary field). Moreover, $\gcd\left(x^{n}-a,x^{m}-b\right)$ has degree $d$ if and only if $a^{m/d}=b^{n/d}$.
@@
Let us now consider $f_{1},f_{2},\ldots,f_{m}$, and put $g:=\gcd(f_{1},f_{2},\ldots,f_{m-1})$. If $g=1$, then $\gcd(g,f_{m})=1$. Else, $g=x^{d^{\prime}}-\displaystyle{\prod_{i=1}^{m-1}}a_{i}^{u_{i}}$, then by applying Lemma 2.8 to $g$ and $f_{m}=x^{n_{m}}-a_{m}$, we obtain
@@ 3 Equivalence of Trinomial Code Families
In this section we study $p^{\ell}$-trinomial codes of length $n=p^{\ell+r}$, where $p$ is the characteristic of the underlying field $\mathbb{F}_{q}$ and $r$ an integer. First, we recall the following lemma, which combines Artin-Schreier’s theorem [Roman1995, Theorem 12.2.1] and [9, Corollary 3.79].
@@
    In particular, if $p=2,$ the polynomial $x^{2}-x-1$ is irreducible over $\mathbb{F}_{2^{2k+1}}$ for any positive integer $k;$ otherwise, it splits in $\mathbb{F}_{2^{2k+1}}.$
  4. 4.
    If $p\neq 2$ then $x^{p}-x-1$ is irreducible over $\mathbb{F}_{2^{2k}}$ for any positive integer $k;$ otherwise, it splits in $\mathbb{F}_{2^{2k}}$.
@@
$(b\alpha^{p^{s+\ell}},a\alpha^{p^{s+k}-p^{\ell}})=(b\alpha^{qp^{\ell}},a\alpha^{(p^{s}-1)p^{\ell}})=(b\alpha^{p^{\ell}},a\alpha^{(q-1)^{p^{\ell}}})=(b\alpha^{p^{\ell}},a)=(1,1).$
@@ 5 Examples of Families of Trinomial Codes
$(a_{0},a_{\ell})\star(\alpha^{3^{\ell+1}},\alpha^{3^{\ell+1}-3^{\ell}})=(a_{0}\alpha^{3^{\ell+1}},a_{\ell}\alpha^{2\cdot 3^{\ell}})=(a_{0}\alpha,a_{\ell})\stackrel{{\scriptstyle!}}{{=}}(1,1).$
@@
$(1,1),(1,\xi),(1,\xi^{2}),(1,\xi^{3}),(1,\xi^{4}),(1,\xi^{5}),(1,\xi^{6}),(1,\xi^{7}).$
It follows that for $a,b\in\mathbb{F}_{9}^{*}$ the $3^{\ell}$-trinomial code family associated with $x^{3^{\ell+2}}-ax^{3^{\ell}}-b,$ is equivalent to a $3^{\ell}$-trinomial code family associated with one of the polynomials $x^{3^{\ell+2}}-x^{3^{\ell}}-\xi^{j},\ j=0,1,\ldots,7.$
@@
    If $k\equiv 0\pmod{3}$: The number of $(27,\ell)$-equivalence classes is $9$. In this case, we consider all pairs $(a,b)$ from $\mathbb{F}_{4}^{*}\times\mathbb{F}_{4}^{*}$.
@@ 6 Equivalence of Polycyclic Codes
    $(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}})\star(\alpha_{i_{0}}^{n-i_{0}},\alpha_{i_{1}}^{n-i_{1}},\ldots,\alpha_{i_{m-1}}^{n-i_{m-1}})=(1,1,\ldots,1).$
@@
    If $\gcd_{0\leq j\leq m-1}\left(\frac{q-1}{\gcd(n-i_{j},q-1)}\right)=1$ then the class of polycyclic codes associated with the polynomial $x^{n}-\displaystyle{\sum_{j=0}^{m-1}}a_{i_{j}}x^{i_{j}}$ is equivalent to the class of polycyclic codes associated with the polynomial $x^{n}-\displaystyle{\sum_{j=0}^{m-1}}\xi^{k_{j}}x^{i_{j}},$ for $k_{j}=0,1,\ldots,\gcd(n-i_{j},q-1),$ where $\xi$ is a primitive element of $\mathbb{F}_{q}$.
  3. 3.
    The number of $\mathbb{F}_{q}$-algebra isometries $\varphi_{\alpha}$ between $\mathbb{F}_{q}[x]/\langle x^{n}-\vec{b}(x)\rangle$ and ${\mathbb{F}_{q}[x]/\langle x^{n}-\vec{a}(x)\rangle}$ is
    $\gcd(n-i_{0},n-i_{1},\ldots,n-i_{m-1},q-1).$

### v2

@@ 2 Preliminaries
    The set $C$ is an ideal of the polynomial ring $\mathbb{F}_{q}[x]/\langle x^{n}-\vec{a}(x)\rangle,$ with $\vec{a}(x)=a_{n-1}x^{n-1}+\ldots+a_{1}x+a_{0}.$
@@
    There is a monic polynomial $g(x)\in\mathbb{F}_{q}[x]$ of least degree which divides $x^{n}-\vec{a}(x)$ and $C=\langle g(x)\rangle.$
@@
    The set $\{g(x),xg(x),\ldots,x^{{n-\deg(g)-1}}g(x)\}$ forms a basis of $C$ and the dimension of $C$ is $n-\deg(g(x)).$
@@
Let $C\subseteq\mathbb{F}_{q}^{n}$ be a non-zero (right) polycyclic code of length $n$ over $\mathbb{F}_{q}$. Then the standard generator of $C$ is the monic polynomial $g(x)$ of least degree in $\mathbb{F}_{q}[x]/\langle f(x)\rangle$ such that $C=\langle g(x)\rangle$. In this paper, we refer to the standard generator of $C$ as ”the generator”.
@@
[15, Lemma 1] Let $m,n\geq 1$ be integers and $a,b\in\mathbb{F}_{q}^{*}$. Then $\gcd\left(x^{n}-a,x^{m}-b\right)$ has degree 0 or $d:=\gcd(m,n)$ (over an arbitrary field). Moreover, $\gcd\left(x^{n}-a,x^{m}-b\right)$ has degree $d$ if and only if $a^{m/d}=b^{n/d}$.
@@
Let us now consider $f_{1},f_{2},\ldots,f_{m}$, and put $g:=\gcd(f_{1},f_{2},\ldots,f_{m-1})$. 5 If $g=1$, then $\gcd(g,f_{m})=1$. Else, $g=x^{d^{\prime}}-\displaystyle{\prod_{i=1}^{m-1}}a_{i}^{u_{i}}$, then by applying Lemma 2.8 to $g$ and $f_{m}=x^{n_{m}}-a_{m}$, we obtain
@@ 3 Equivalence of Trinomial Code Families
Let $\ell$ be an integer such that $0<\ell<n$, and $(a_{0},a_{\ell})$ be an element of $\mathbb{F}_{q}^{*}\times\mathbb{F}_{q}^{*}$. Then the following statements are equivalent.
@@
In this section we study $p^{\ell}$-trinomial codes of length $n=p^{\ell+r}$, where $p$ is the characteristic of the underlying field $\mathbb{F}_{q}$ and $r$ an integer. First, we recall the following lemma, which combines Artin-Schreier’s theorem [16, Theorem 12.2.1] and [9, Corollary 3.79].
@@
    In particular, if $p=2,$ the polynomial $x^{2}-x-1$ is irreducible over $\mathbb{F}_{2^{2k+1}}$ for any positive integer $k;$ otherwise, it splits in $\mathbb{F}_{2^{2k}}.$
@@ 8 Now we can factor $x^{p^{\ell+1}}-ax^{p^{\ell}}-b$ as follows:
$(b\alpha^{p^{s+\ell}},a\alpha^{p^{s+\ell}-p^{\ell}})=(b\alpha^{qp^{\ell}},a\alpha^{(p^{s}-1)p^{\ell}})=(b\alpha^{p^{\ell}},a\alpha^{(q-1)^{p^{\ell}}})=(b\alpha^{p^{\ell}},a)=(1,1).$
@@ 5 Examples of Families of Trinomial Codes
$(a_{0},a_{\ell})\star(\alpha^{3^{\ell+1}},\alpha^{3^{\ell+1}-3^{\ell}})=(a_{0}\alpha^{3^{\ell+1}},a_{\ell}\alpha^{2\cdot 3^{\ell}})=(a_{0}\alpha,a_{\ell})=(1,1).$
@@
$(1,1),(\xi,1),(\xi^{2},1),(\xi^{3},1),(\xi^{4},1),(\xi^{5},1),(\xi^{6},1),(\xi^{7},1).$
It follows that for $a,b\in\mathbb{F}_{9}^{*}$ the $3^{\ell}$-trinomial code family associated with $x^{3^{\ell+2}}-ax^{3^{\ell}}-b,$ is equivalent to a $3^{\ell}$-trinomial code family associated with one of the polynomials $x^{3^{\ell+2}}-\xi^{j}x^{3^{\ell}}-1,\ j=0,1,\ldots,7.$
@@
    If $\ell\equiv 0\pmod{3}$: The number of $(27,\ell)$-equivalence classes is $9$. In this case, we consider all pairs $(a,b)$ from $\mathbb{F}_{4}^{*}\times\mathbb{F}_{4}^{*}$.
@@ 6 Equivalence of Polycyclic Codes
    $(a_{i_{0}},a_{i_{1}},\ldots,a_{i_{m-1}})\star(\alpha^{n-i_{0}},\alpha^{n-i_{1}},\ldots,\alpha^{n-i_{m-1}})=(1,1,\ldots,1).$
@@
    Let $d=\gcd_{0\leq j\leq m-1}\left(\frac{q-1}{\gcd(n-i_{j},q-1)}\right),$ then the class of polycyclic codes associated with the polynomial $x^{n}-\displaystyle{\sum_{j=0}^{m-1}}a_{i_{j}}x^{i_{j}}$ is equivalent to the class of polycyclic codes associated with the polynomial $x^{n}-\displaystyle{\sum_{j=0}^{m-1}}\xi^{k_{j}}x^{i_{j}},$ for $k_{j}=0,1,\ldots,\gcd(n-i_{j},q-1)-1,$ where $\xi$ is a primitive element of $\mathbb{F}_{q}$.
Proof.
Proof.
  1. 1.
    Follows from the fifth assertion of Theorem of 6.4.
  2. 2.
    Let $d=\gcd_{0\leq j\leq m-1}\left(\frac{q-1}{\gcd(n-i_{j},q-1)}\right).$
    • (a)
      If $d=1$, then the cyclic group $H$ generated by $(\xi^{n-{i_{0}}},\xi^{n-{i_{1}}},\ldots,\xi^{n-{i_{m-1}}})$ is isomorphic to the group $\langle\xi^{n-{i_{0}}}\rangle\times\langle\xi^{n-{i_{1}}}\rangle\times\ldots\times\langle\xi^{n-{i_{m-1}}}\rangle$ and has order $\frac{(q-1)^{m}}{d_{0}d_{1}\ldots d_{m-1}},$ with $d_{j}=\gcd(n-i_{j},q-1)$ for $j=0,\ldots,m-1$. By Theorem 6.4, the number of $n$-equivalence classes is $d_{0}d_{1}\ldots d_{m-1}$. Therefore, we can partition $(\mathbb{F}_{q}^{*})^{m}$ as
      $(\mathbb{F}_{q}^{*})^{m}=\bigcup_{k_{0}=0}^{d_{0}-1}\bigcup_{k_{1}=0}^{d_{1}-1}\ldots\bigcup_{k_{m-1}=0}^{d_{m-1}-1}(\xi^{k_{0}},\xi^{k_{1}},\ldots,\xi^{k_{m-1}})H.$
      So the result holds.
    • (b)
      If $d\neq 1,$ the number of $(n,\ell)$-equivalence classes is $dd_{0}d_{1}\ldots d_{m-1},$ and so we partition $(\mathbb{F}_{q}^{*})^{m}$ as
      $(\mathbb{F}_{q}^{*})^{m}=\bigcup_{h=0}^{d-1}\bigcup_{k_{0}=0}^{d_{0}-1}\bigcup_{k_{1}=0}^{d_{1}-1}\ldots\bigcup_{k_{m-1}=0}^{d_{m-1}-1}(\xi^{k_{0}+hn},\xi^{k_{1}},\ldots,\xi^{k_{m-1}})H.$
      which implies the result.